Find the length of a ladder leaning against a building if the top of the ladder touches the building at a height of 12 feet. Also the length of the ladder is 4 feet more than its distance from the base of the building.
n=4n+12

Answer by checkley77(12844)

12^2+x^2=(x+4)^2
144+x^2=x^2+8x+16
8x=144-16
8x=128
x=128/8
x=16 ft. is the distance from the building to the foot of the ladder
16=4=20 ft. is the length of the ladder.
Proof:
12^2=16^2=20^2
144+1256=400
400=400